# Fed-PISA: Federated Voice Cloning via Personalized Identity-Style Adaptation

### v1

@@ 3.1.1 Datasets
Table 1: Statistics of datasets used in our study, including the number of heterogeneous styles, speakers, and the train/val/test splits for both neutral and expressive subsets.
@@ 3.1.3 Training Details
The federated simulation comprises 60 clients, where each speaker from the datasets is treated as an independent client. The entire training process spans 50 communication rounds, with a client participation rate of 20% per round. For both the private ID-LoRA and the federated Style-LoRA, we set the rank to 8 and the scaling factor to $16$. Local adaptation is performed using the AdamW optimizer with a batch size of $16$. We employ a learning rate of $2\times 10^{-5}$ with a cosine decay schedule and a warmup ratio of $0.1$. For main experiments on each client, the timbre cloning phase runs for $80$ steps, followed by the stylization phase, which runs for $20$ steps. Experiments were conducted on four NVIDIA V100 GPUs.
@@ 5 Compliance with Ethical Standards
Compliance with Ethical Standards. This study was performed in line with the principles of the Declaration of Helsinki. Approval was granted by the Ethics Committee of the Institute of Computing Technology, Chinese Academy of Sciences (ICT, CAS). All participants in the blind listening tests (n = 22) provided informed consent; only anonymized ratings were collected. We additionally used open-access human-speech datasets (ESD, EmoV-DB, RAVDESS, CREMA-D) strictly in accordance with their licenses; no new recordings were collected from these corpora.

### v2

@@ 3.1.1 Datasets
Table 1: Statistics of datasets used in our study, including the number of heterogeneous styles, speakers, and the train/val/test sample splits for both neutral and expressive subsets.
@@ 3.1.3 Training Details
Each speaker from the datasets is treated as an independent client. The entire training process spans 50 communication rounds, with a client participation rate of 20% per round. For both the private ID-LoRA and the federated Style-LoRA, we set the rank to 8 and the scaling factor to $16$. Local adaptation is performed using the AdamW optimizer with a batch size of $16$. We employ a learning rate of $2\times 10^{-5}$ with a cosine decay schedule and a warmup ratio of $0.1$. For main experiments on each client, the timbre cloning phase runs for $80$ steps, followed by the stylization phase, which runs for $20$ steps. Experiments were conducted on four NVIDIA V100 GPUs.
@@ 5 Compliance with Ethical Standards
Compliance with Ethical Standards. This study was performed in line with the principles of the Declaration of Helsinki. Approval was granted by the Ethics Committee of the Institute of Computing Technology, Chinese Academy of Sciences (ICT, CAS). All participants in the blind listening tests provided informed consent. Only anonymized ratings were collected. We additionally used open-access human-speech datasets strictly in accordance with their licenses. No new recordings were collected from these corpora.